Verify each indentity.
1. cscxsecx=cotx+tanx

Answer by mananth(16946)

cscxsecx=cotx+tanx
=1/six * 1/cosx
=1/sinxcosx
= sin^2x+cos^2/sinxcosx
= sin^2/sinxcosx + cos^2x/sinxcosx
=sinx/cosx+cosx/sinx
=tanx+cotx